United States Environmental Protection Agency
Region X
POLLUTION REPORT



Date:
Monday, July 14, 2003
From:
Michael Boykin

To:
Ryan Whitchurch, E & E START

Subject: 

Final POLREP
Hwy 42 Waste Oil Spill
Milepost 1.2 Hwy 42, Coos Bay, OR
Latitude: 43.2931000
Longitude: -124.2419000


POLREP No.:
3
Site #:
Z0A6
Reporting Period:
06/27-30/03
D.O. #:
N/A
Start Date:
6/23/2003
Response Authority:
OPA
Mob Date:
6/23/2003
Response Type:
Demob Date:
 
NPL Status:
Completion Date:
 
Incident Category:
CERCLIS ID #:
N/A
Contract #
N/A
RCRIS ID #:
Reimbursable Account #
FPN#
E03015
 

Site Description

For Site Description please refer to Initial POLREP.


Current Activities

EPA OSC demobed from site to Seattle on Thursday, June 26, 2003 and designated USCG PST member as the FOSC representative in Unified Command.

Friday, June 27, 2003: Oregon State OSC (1), USCG PST (1).
1. Excavation of contaminated soil in the median strip proceeded to within limits established by ODEQ and ODOT:  no closer to buried fiber optic cable than 6-inches (verified by on-site Verizon representative), no closer to Hwy 42 eastbound lane than directed by ODOT on-site representative (excavation into shoulder and edge of road was performed).
2. Foss continued Hwy 42 "West Wall" median surface soil excavation under ODOT supervision.
3. ODOT re-paved/repaired damaged eastbound lanes of Hwy 42.
4. Pacific Rail authorized Foss to conduct push-probe operations to within 10-feet of rail line adjacent to slough.
5. SOSC and PBS Environmental representative collected 33 confirmation samples from Hwy 42 median strip excavation area (-60-feet to +200-feet from spill, along bottom and sides).  All confirmation samples to be analyzed for NWTPH-Gx, NWTPH-Dx, and selected samples for VOCs (8260b) and semi-volatiles (8270c).  Additional median confirmation samples are also anticipated.
6. ODOT completed temporary paving repairs to test pit area of fast lane of westbound Hwy 42.
7. SOSC requested specific analyses of one background soil, one product/soil, and three surface-water samples (8260b, 8270c, PCB, Metals) collected by EPA/START.
8. Coffin Butte Landfill informed Foss that TCLP analyses are required prior to acceptance of contaminated soils.
9. An estimated 130 cubic yards of soil excavated during June 27, 2003.
10. Eight ¾-inch monitoring wells installed using push-probe on Isthmus Slough side of Hwy 42.
11. Skirt containment boom removed from Isthmus Slough by Foss.
12. Double layered sorbent boom placed around culvert that discharges into Isthmus Slough.
13. Preliminary results for waste oil in soil indicate presence of gasoline, diesel, tetrachloroethylene (PCE), and styrene.

Saturday, June 28, 2003 SOSC (1), USCG PST (1)
1. Johnson Rock (subcontracted to Foss) begin backfill operations of median strip excavation.
2. Foss makes available full Site Specific Health and Safety Plan; previously an emergency tailgate-version plan was available.
3. Foss notifies Pacific Railroad that push-probe activities near rail are completed.
4. Two push-probe ¾-inch monitoring wells installed upgradient, along Hwy 42 eastbound median and shoulder.  All wells are pre-packed construction, and were all installed by Pacific Northwest Drilling contracted to Foss with ODEQ geologist oversight.
5. Six power-probe drill site locations (2 in fast lane of eastbound Hwy 42 and 4 in fast lane of westbound Hwy 42) completed for soil sampling beneath road surface.  No elevated PID readings observed and all samples to be archived.
6. USCG PST demobs from site at 1000 hours to return to Portland.
7. An estimated 90 cubic yards of soil excavated on June 28, 2003.
8. Contaminated soil consolidated at staging area, and soil stockpiles covered with plastic sheeting.
9. All traffic control measures removed and operations completed by early evening.
10. A total of 298 feet of road surface repaired.

Sunday, June 29, 2003  SOSC (1)
1. Survey of soil sample and monitoring well locations conducted.
2. The 10 monitoring wells were sampled by Colin Polk of PBS (subcontractor to Foss).  Reportedly, no odors, sheen, or elevated PID measurements were observed during sampling of groundwater monitoring wells.
3. No soil excavated today and stockpiled soil remains staged and covered.
4. No sheen noted in Isthmus Slough or Wall Creek.  Absorbent boom to be maintained for possible sheen during precipitation.
5. Total soil excavated during this response cleanup is estimated to be 950 cubic yards.
6. Stockpiled soil remains staged and covered pending disposal scheduled for 7/10/03 and 7/11/03.
7. Verizon fiber optic cable not damaged during site cleanup activities.
8. SOSC indicates that the emergency response cleanup action is complete.
9. SOSC and Foss and subcontractors demobilize from site.


Planned Removal Actions

Foss to arrange for transportation and disposal of stockpiled soils to occur July 10-11, 2003.

No further removal actions are anticipated at this time.


Next Steps

Site to be transferred to Oregon Department of Environmental Quality Cleanup Program for assessment, groundwater monitoring, and long-term clean-up.

SOSC to arrange for DEQ personnel to monitor Isthmus Slough and Wall Creek for sheen/releases during the next significant precipitation event.


Disposition of Wastes

Date            Amount of Wastes        Disposition
June 24, 2003   85  cu. yds. soil       Held in Staging Area 1
June 25, 2003  425 cu. yds. soil       Held in Staging Area 1
June 26, 2003  220 cu. yds. soil       Held in Staging Area 1
June 27, 2003  130 cu. yds. soil       Held in Staging Area 1
June 28, 2003   90  cu. yds. soil      Held in Staging Area 1
Total          950 cu. yds. soil



response.epa.gov/Hwy42WasteOilSpill